[figure: 9.7 cm]

GIANT MAGNETOOPTICS OF BISMUTH-SUBSTITUTED IRON

GARNETS: A THEORETICAL ANALYSIS

A.V. Zenkov and A.S. Moskvin Dept. of Theoretical Physics Ural State A.M. Gorky University Lenin Ave., 51, 620083 Ekaterinburg, Russia

In the framework of the charge-transfer transitions concept, we suggest a semi-quantitative model which explains the strong increase of the circular magnetooptics of iron garnets R_{3}Fe_{5}O_{12} when doped with Bi^{3+} or Pb^{2+} ions. The covalent admixture of Bi^{3+}, Pb^{2+} 6p-orbitals (with giant one-electron spin-orbit coupling constant) to oxygen 2p-states causes the growth of the oxygen contribution to the spin-orbit coupling constant of (FeO_{6})^{9−}, (FeO_{4})^{5−} complexes (main magnetooptically active centers of iron garnets). The enhanced spin-orbit interaction on the oxygen is not the only effect of substitution; it also gives rise to the effective anisotropic tensor addend to the spin-orbit interaction and the circular magnetooptics of iron garnets as well. Estimates of various contributions to the iron garnets magnetooptics are given. The computer simulation of magnetooptic spectra of Y_{3−x}Bi_{x}Fe_{5}O_{12} under the inhomogeneous bismuth distribution is made. Analysis of experimental data on garnets magnetooptics speaks in favor of our theoretical model.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : A.V. Zenkov

Address for correspondence : Kuibyshev St., 88-43, Ekaterinburg 620100, RUSSIA

Email address : andreas@r66.ru